find function f such that g(x)= 1/x-5
h=g o f, and h(x)=1/(x^2-5)
Thank you!

Answer by stanbon(75887)

find function f such that g(x)= 1/x-5
h=g o f, and h(x)=1/(x^2-5)
------------------------------
h(x) = gof(x) = g[f(x)] = 1/[f(x)-5)]
----
But h(x) = 1/(x^2-5)
----------
1/[f(x)-5] = 1/[x^2-5]
f(x)-5 = x^2-5
f(x) = x^2
================
Cheers,
Stan H.